# Explaining Object Detectors via Collective Contribution of Pixels

### v2

@@ Abstract
Visual explanations for object detectors are crucial for enhancing their reliability. Object detectors identify and localize instances by assessing multiple visual features collectively. When generating explanations, overlooking these collective influences in detections may lead to missing compositional cues or capturing spurious correlations. However, existing methods typically focus solely on individual pixel contributions, neglecting the collective contribution of multiple pixels. To address this limitation, we propose a game-theoretic method based on Shapley values and interactions to explicitly capture both individual and collective pixel contributions. Our method provides explanations for both bounding box localization and class determination, highlighting regions crucial for detection. Extensive experiments demonstrate that the proposed method identifies important regions more accurately than state-of-the-art methods. The code will be publicly available soon.
@@ 1 Introduction
Figure 1: Generated heat maps by ODAM, SSGrad-CAM++, and VX-CODE. Each heat map shows the insertion AUC (Ins AUC), which measures faithfulness (higher is better). While previous methods highlight dominant features (e.g., hand or surfboard), VX-CODE captures collective contributions across multiple features (e.g., leg+head or surfboard+sea) through interactions.
@@ 2 Related Work
Gradient-based methods compute instance-specific importance by propagating gradients through the network [19, 12, 65]. ODAM [65] multiplies feature and gradient maps to estimate importance. CAM-based methods, like SSGrad-CAM [60] and SSGrad-CAM++ [59], extend Grad-CAM [44] by incorporating spatial information. Perturbation-based methods measure changes in the detector’s output when parts of the input images are masked [39, 55, 8]. D-RISE [39], an extension of RISE [38], masks inputs randomly, while BSED [27] and FSOD [28] use Shapley values to estimate contributions.
@@ 3 Preliminaries
Importantly, when two patches $i$ and $j$ encode nearly the same information and the merged player adds no more than either one alone, i.e., $\phi(S_{ij}\mid N^{\prime})\approx\phi(i\mid N\setminus\{j\})\approx\phi(j\mid N\setminus\{i\})=:m>0$, the interaction becomes negative (if each ). This characteristic is crucial for identifying a minimal, non-redundant set of patches that well explain object detectors. If one selects only patches with high Shapley values as in prior study, patches containing assistive information (e.g., background) may be overlooked, as they typically contribute less individually compared to those containing primary information (e.g., the body of objects).
@@
where $\mathrm{IoU}(L^{x},L)$ measures localization and the cosine term measures classification agreement. Note that in general, $(L^{x},P^{x})\notin\mathcal{D}(x_{S})$ unless $S=N$. See App. A for its validation, and App. D.10 for a further analysis of the effect of the reward formulation on the resulting explanations.
@@ 4 Approach
Explaining biased model. We train a biased model on a dataset where 70% of the images contain a square marker placed outside the upper-right corner of the bounding box. As a result, the trained model relies on both objects and the marker for detection. In Fig. 2, VX-CODE, which considers collective contributions, clearly identifies both cues. Previous methods highlight only the square marker (Grad-CAM++ and D-RISE) or almost the entire bounding box (SSGrad-CAM++). The AUC score quantitatively confirm these observations. See App. D.7 for additional results.
@@
Explaining failure cases. Figure 3 presents two failure cases of a standard detection model. The first case is a misclassification of a dog as a bird due to the presence of sky, and the second is a mislocalization of the bicycle caused by the human. VX-CODE, considering collective contributions, correctly identifies the causes of these misdetections: both the sky and the human’s body (e.g., face and legs) in the second. ODAM and SSGrad-CAM++ highlight only the body or human-related features, whereas D-RISE produces noisy explanations. In both cases, the AUC score quantitatively supports these findings. See Sec. 5.3 and App. D.12 for additiona results.
@@ 4.1 Proposed method
To this end, we propose VX-CODE, an efficient heuristic to address Problem 1. Figure 4 provides an overview of the computation pipeline. Here, we focus on the insertion setup (patch insertion); the deletion setup (patch deletion) is methodologically similar and deferred to App. C.4.
@@
The self-context values were introduced in [51] as a conceptual analogy of Shapley values and interactions. In this work, we reveal the connection between our framework and the original concepts. We also introduce the $\pi^{*}$-index, which provides a perspective from sequential coalitional games and show that this index satisfies the essential axioms of coalitional games (see Sec. 4.3 and App. C.8). Next, we define self-context interaction based on Eq. (2), with a slight generalization to handle two or more players:
@@
Initial step (k=1). The set of patches $B_{1}\ subsetN$ that maximizes $f(B_{1})$ is selected.
@@ 4.3 Analysis
The self-context Shapley values and interactions are inspired by the original indices, and do not satisfy some axioms of coalitional games, e.g., efficiency. Although this does not affect the practical utility, we introduce the $\pi^{*}$-index, which satisfies the axioms (cf. App. C.8) and provides a perspective on our strategy as a sequential coalitional game. Recall $N=\{1,\ldots,n\}$ is the set of player IDs. Let $\Pi$ be the set of all permutations over $N$, where $\pi\in\Pi$ permutes a list $1,\ldots,n]$ to $\pi(1),\ldots,\pi(n)]$.
@@
Choice of $r$. Our framework considers both Shapley values and interactions, and increasing $r$ enables evaluation over more patches (see App. C.3 for examples with $r=1$ and $r=2$). While $r=1$ already works well (see Sec. 5.1), larger $r$ further improves the identification of important patches. In greedy selection, each step affects the next; thus, failing to select important patches early can lead to an uninformative search. When $r=1$, limited early information may hinder the identification of important patches, and we observe that some cases fail as a result. Increasing $r$ mitigates this by allowing consideration of various subset sizes and patch interactions (see Sec. 5.2 and Fig. 7).
@@ 5 Experiments
We compare VX-CODE with Grad-CAM [44], Grad-CAM++ [7], D-RISE [39], SSGrad-CAM [60], ODAM [65], and SSGrad-CAM++ [59]. For D-RISE, we use 5,000 masks at $16\times 16$ resolution. The same smoothing function is applied to SSGrad-CAM, ODAM, and SSGrad-CAM++. For evaluation, we use MS-COCO [29] and PASCAL VOC [13] datasets, which provide diverse scenes and object categories for robust comparison. We apply each method to two object detectors: DETR [6] and Faster R-CNN [42]. [^3] For PASCAL VOC, both models are trained with ImageNet pre-trained weights. For Faster R-CNN on MS-COCO, we use the pre-trained models from [58]. For DETR on MS-COCO, we use the model from https://github.com/facebookresearch/detr.git. We use ResNet-50 [20] as the backbone for DETR, and ResNet-50 and FPN [30] for Faster R-CNN. All implementations are based on detectron2 [58], a PyTorch-based library [37]. In VX-CODE, heat maps are generated from selected patches, as described in App. C.6. Implementation details are in App. C.7. To simulate patch removal in VX-CODE, we apply local zero masking [39]. We also evaluate blurring [14, 46] as an alternative masking. Additional results are provided in App. D.1.
@@ 5.1 Faithfulness of identified regions
We compare our method with Monte Carlo-based Shapley values (see App. D.2). We also evaluate class-level and box-level explanations separately (see App. D.8 and D.9) and localization performance using the pointing game [62] (see App. D.4).
@@ 5.2 Qualitative analysis of visualizations
Figure 6 shows explanations from each method for DETR detections. In the first row of Fig. 6 (a), VX-CODE identifies multiple car features, including the body, windshield, and shovel, which jointly contribute to the detection. In contrast, the other methods either highlight irrelevant regions or focus only on the shovel. In the second row of Fig. 6 (a), VX-CODE with patch insertion captures both the surfboard and the surrounding sea surface, indicating that VX-CODE identifies background context as contributing to the detection. In contrast, the other methods focus only on the surfboard and are less sensitive to the background. VX-CODE with patch deletion focuses solely on the surfboard. Patch deletion prioritizes removing instance features, which produce the largest drop in detection confidence. As a result, the resulting explanations are object-centric and can improve object localization. We provide additional results in App. D.5, along with sanity checks, described in App. D.3.
@@
Table 2: Results of the deletion (Del), insertion (Ins), and overall (OA) for failure cases of mislocalization and misclassification from DETR on MS-COCO. Each result is computed over 200 detection instances for each failure case.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP23K24914, JST PRESTO Grant Number JPMJPR24K4, and JST BOOST Program Grant Number JPMJBY24C6.
@@ Appendix B Visual explanations for classification
Methods that consider interactions [17] in addition to Shapley values have also been proposed [4, 51]. PredDIFF [4] is a method that measures changes in predictions by marginalizing features from a probabilistic perspective. This paper clarifies the relationship between PredDIFF and Shapley values, and introduces a new measurement method that takes interactions into account. MoXI [51] utilizes a technique to approximate Shapley values and interactions to identify important regions for classification models. While MoXI considers only pairwise patch interactions and their Shapley values (equivalent to $r=1$ in our method), our method considers interactions with arbitrary patches and their Shapley values. As described in Sec. 5.2, in explanations for object detectors, some cases fail to identify important patches when using $r=1$. The proposed method addresses this limitation and provides more faithful explanations for object detectors. Additionally, we introduce $\pi^{*}$-index that clarifies the connection between its underlying concept and sequential coalitional games (see Sec. 4.3), and prove that this index satisfies the essential axioms of coalitional games (see App. C.8).
@@ C.2 Details of the algorithm
C.3 Specific example of patch insertion
@@ C.4 Patch deletion
C.5 Specific example of patch deletion
We provide specific examples for cases with $r=1$ and $r=2$ in patch deletion described in App. C.4.
C.5.1 Case with r=1
@@ C.5.2 Case with r=2
C.6 Heat maps generation from identified patches
@@ C.7 Implementation details
Figure H: Comparison results of the two filling methods, zero filling and blurred pixel values, using (a) the insertion metric and (b) the deletion metric. The similarity score is computed using Eq. (3).
Figure I: The comparison of insertion and deletion metrics between VX-CODE and Shapley values estimated via Monte Carlo sampling.
Figure J: Comparison of heat maps generated by VX-CODE with patch insertion for the original and weight-randomized models.
@@
C.8 Axiomatic properties of the $\pi^{*}$-index
@@
The $\psi^{\mathrm{es}}$ measures the contribution of each player by the decrease in reward upon their removal. It is worth noting that the greedy patch deletion process described in App. C.4 can be interpreted as the player elimination setting in a sequential coalitional game. As shown in Eq. 62, at each step the algorithm selects the set of patches whose removal produces the largest decrease in the reward, which is equivalent to selecting the set with the largest full-context Shapley value. In this sense, the patch deletion process procedure induces a specific permutation $\hat{\pi}$ that defines a sequential coalitional game in the player elimination setting.
@@ Appendix D Additional experimental results
Figure K: Additional results comparing the visual explanations generated by each method for detections from DETR. In VX-CODE, we set $r=1$. (a) Heat maps generated by each method. (b) Patches identified from VX-CODE with patch insertion (left) and patch deletion (right). Each heat map corresponds to those shown in (a).
Figure L: Additional results comparing the visual explanations generated by each method for detections from Faster R-CNN. In VX-CODE, we set $r=1$. (a) Heat maps generated by each method. (b) Patches identified from VX-CODE with patch insertion (left) and patch deletion (right). Each heat map corresponds to those shown in (a).
Figure M: Additional results comparing the identified patches and generated heat maps by VX-CODE when changing $r$. (a) Results for patch insertion. (b) Results for patch deletion.
Figure N: (a) Example of synthetic data: A square marker is placed in the upper-right corner outside the bounding box. These data include annotations for the class and bounding box. (b) Example of data without annotations: These data lack annotations and are treated as background during training.
@@ D.1 Comparison of filling methods
As described in Sec. 5, we fill the masked regions with zeros in the proposed method. Here, we compare two filling methods, zero filling and blurred pixel values, using the insertion and deletion metrics. Figure H shows the results. We observe that both methods produce nearly identical results. These findings suggest that the choice of filling method has no significant impact on the proposed method.
@@ D.2 Comparison of Monte-Carlo based Shapley values
We compare the proposed method with Shapley values estimated via Monte Carlo sampling. Figure I shows the results of the insertion and deletion metrics, computed for 100 instances detected by DETR on MS-COCO. The number of Monte Carlo samples is set to 200, following prior studies, which already incurs a higher computational cost than our method with $r=1$. As shown in Fig. I, the proposed method with $r=1$ already outperforms the Shapley values estimated by Monte Carlo sampling.

### v4

@@ Abstract
Visual explanations for object detectors are crucial for enhancing their reliability. Object detectors identify and localize instances by assessing multiple visual features collectively. When generating explanations, overlooking these collective influences in detections may lead to missing compositional cues or capturing spurious correlations. However, existing methods typically focus solely on individual pixel contributions, neglecting the collective contribution of multiple pixels. To address this limitation, we propose a game-theoretic method based on Shapley values and interactions to explicitly capture both individual and collective pixel contributions. Our method provides explanations for both bounding box localization and class determination, highlighting regions crucial for detection. Extensive experiments demonstrate that the proposed method identifies important regions more accurately than state-of-the-art methods. The code is available at https://github.com/tttt-0814/VX-CODE.
@@ 1 Introduction
Figure 1: Generated heat maps by D-RISE, ODAM, SSGrad-CAM++, and VX-CODE. Each heat map shows the insertion AUC (Ins AUC), which measures faithfulness (higher is better). While previous methods highlight dominant features (e.g., hand or surfboard), VX-CODE captures collective contributions across multiple features (e.g., leg+head or surfboard+sea) through interactions.
@@ 2 Related Work
Gradient-based methods compute instance-specific importance by propagating gradients through the network [19, 12, 66]. ODAM [66] multiplies feature and gradient maps to estimate importance. CAM-based methods, like SSGrad-CAM [61] and SSGrad-CAM++ [60], extend Grad-CAM [45] by incorporating spatial information. Perturbation-based methods measure changes in the detector’s output when parts of the input images are masked [40, 56, 8]. D-RISE [40], an extension of RISE [39], masks inputs randomly, while BSED [27] and FSOD [28] use Shapley values to estimate per-pixel or per-region importance.
@@ 3 Preliminaries
Importantly, when two patches $i$ and $j$ encode nearly the same information and the merged player adds no more than either one alone, i.e., $\phi(S_{ij}\mid N^{\prime})\approx\phi(i\mid N\setminus\{j\})\approx\phi(j\mid N\setminus\{i\})=:m>0$, the interaction becomes negative. This characteristic is crucial for identifying a minimal, non-redundant set of patches that well explain object detectors. If one selects only patches with high Shapley values as in prior study, patches containing assistive information (e.g., background) may be overlooked, as they typically contribute less individually compared to those containing primary information (e.g., the body of objects).
@@
where $\mathrm{IoU}(L^{x},L)$ measures localization and the cosine term measures classification agreement. Note that in general, $(L^{x},P^{x})\notin\mathcal{D}(x_{S})$ unless $S=N$. We provide a conceptual illustration of this computation in App. C.3, validate it in App. A, and further analyze the effect of the reward formulation on the resulting explanations in App. D.10.
@@ 4 Approach
Explaining biased model. We train a biased model on a dataset where 70% of the images contain a class-agnostic square marker placed outside the upper-right corner of the bounding box. As a result, the trained model relies on both objects and the marker for detection. In Fig. 2, VX-CODE, which considers collective contributions, clearly identifies both cues. Previous methods highlight only the square marker (Grad-CAM++ and D-RISE) or almost the entire bounding box (SSGrad-CAM++). The AUC score quantitatively confirm these observations. See App. D.7 for additional results.
@@
Explaining failure cases. Figure 3 presents two failure cases of a standard detection model. The first case is a misclassification of a dog as a bird due to the presence of sky, and the second is a mislocalization of the bicycle caused by the human. VX-CODE, considering collective contributions, correctly identifies the causes of these misdetections: both the sky and the human’s body (e.g., face and legs) in the second. ODAM and SSGrad-CAM++ highlight only the body or human-related features, whereas D-RISE produces noisy explanations. In both cases, the AUC score quantitatively supports these findings. See Sec. 5.3 and App. D.12 for additional results.
@@ 4.1 Proposed method
To this end, we propose VX-CODE, an efficient heuristic to address Problem 1. Figure 4 provides an overview of the computation pipeline. Here, we focus on the insertion setup (patch insertion); the deletion setup (patch deletion) is methodologically similar and deferred to App. C.5.
@@
The self-context values were introduced in [52] as a conceptual analogy of Shapley values and interactions. In this work, we reveal the connection between our framework and the original concepts. We also introduce the $\pi^{*}$-index, which provides a perspective from sequential coalitional games and show that this index satisfies the essential axioms of coalitional games (see Sec. 4.3 and App. C.9). Next, we define self-context interaction based on Eq. (2), with a slight generalization to handle two or more players:
@@
Initial step (k=1). The set of patches $B_{1}\subset N$ that maximizes $f(B_{1})$ is selected.
@@ 4.3 Analysis
The self-context Shapley values and interactions are inspired by the original indices, and do not satisfy some axioms of coalitional games, e.g., efficiency. Although this does not affect the practical utility, we introduce the $\pi^{*}$-index, which satisfies the axioms (cf. App. C.9) and provides a perspective on our strategy as a sequential coalitional game. Recall $N=\{1,\ldots,n\}$ is the set of player IDs. Let $\Pi$ be the set of all permutations over $N$, where $\pi\in\Pi$ permutes a list $1,\ldots,n]$ to $\pi(1),\ldots,\pi(n)]$.
@@
Choice of $r$. Our framework considers both Shapley values and interactions, and increasing $r$ enables evaluation over more patches (see App. C.4 for examples with $r=1$ and $r=2$). While $r=1$ already works well (see Sec. 5.1), larger $r$ further improves the identification of important patches. In greedy selection, each step affects the next; thus, failing to select important patches early can lead to an uninformative search. When $r=1$, limited early information may hinder the identification of important patches, and we observe that some cases fail as a result. Increasing $r$ mitigates this by allowing consideration of various subset sizes and patch interactions (see Sec. 5.2 and Fig. 7).
@@ 5 Experiments
We compare VX-CODE with Grad-CAM [45], Grad-CAM++ [7], D-RISE [40], SSGrad-CAM [61], ODAM [66], and SSGrad-CAM++ [60]. For D-RISE, we use 5,000 masks at $16\times 16$ resolution. The same smoothing function is applied to SSGrad-CAM, ODAM, and SSGrad-CAM++. For evaluation, we use MS-COCO [29] and PASCAL VOC [13] datasets, which provide diverse scenes and object categories for robust comparison. We apply each method to two object detectors: DETR [6] and Faster R-CNN [43]. [^3] For PASCAL VOC, both models are trained with ImageNet pre-trained weights. For Faster R-CNN on MS-COCO, we use the pre-trained models from [59]. For DETR on MS-COCO, we use the model from https://github.com/facebookresearch/detr.git. We use ResNet-50 [20] as the backbone for DETR, and ResNet-50 and FPN [30] for Faster R-CNN. All implementations are based on detectron2 [59], a PyTorch-based library [38]. In VX-CODE, heat maps are generated from selected patches, as described in App. C.7. Implementation details are in App. C.8. To simulate patch removal in VX-CODE, we apply local zero masking [40]. We also evaluate blurring [14, 47] as an alternative masking. Additional results are provided in App. D.1.
@@ 5.1 Faithfulness of identified regions
We evaluate VX-CODE on the one-stage detector RetinaNet [31] (App. D.15), compare it with Monte Carlo-based Shapley values (App. D.2), and further analyze collective effects from the interaction perspective (App. D.16). We also evaluate class-level and box-level explanations separately (App. D.8 and D.9), and assess localization performance using the pointing game [63] (App. D.4).
@@ 5.2 Qualitative analysis of visualizations
Figure 6 shows explanations from each method for DETR detections. In the first row of Fig. 6 (a), VX-CODE identifies multiple car features, including the body, windshield, and shovel, which jointly contribute to the detection. In contrast, the other methods either highlight irrelevant regions or focus only on the shovel. In the second row of Fig. 6 (a), VX-CODE with patch insertion captures both the surfboard and the surrounding sea surface, indicating that VX-CODE identifies background context as contributing to the detection. In contrast, the other methods focus only on the surfboard and are less sensitive to the background. VX-CODE with patch deletion focuses solely on the surfboard. Patch deletion prioritizes removing instance features, which produce the largest drop in detection confidence. As a result, the resulting explanations are object-centric and can improve object localization. We provide additional results in Apps. D.5 and D.14, along with sanity checks (see App. D.3).
@@
Table 2: Results of the deletion (Del) and insertion (Ins) for failure cases of mislocalization and misclassification from DETR on MS-COCO. Each result is computed over 200 detection instances for each failure case.
@@ Acknowledgments
This work was supported by JSPS KAKENHI Grant Number JP23K24914, JST PRESTO Grant Number JPMJPR24K4, JST BOOST Program Grant Number JPMJBY24C6, and ROIS NII Open Collaborative Research 261S07-24168.
@@ Appendix B Visual explanations for classification
Methods that consider interactions [17] in addition to Shapley values have also been proposed [4, 52]. PredDIFF [4] is a method that measures changes in predictions by marginalizing features from a probabilistic perspective. This paper clarifies the relationship between PredDIFF and Shapley values, and introduces a new measurement method that takes interactions into account. MoXI [52] utilizes a technique to approximate Shapley values and interactions to identify important regions for classification models. While MoXI considers only pairwise patch interactions and their Shapley values (equivalent to $r=1$ in our method), our method considers interactions with arbitrary patches and their Shapley values. As described in Sec. 5.2, in explanations for object detectors, some cases fail to identify important patches when using $r=1$. The proposed method addresses this limitation and provides more faithful explanations for object detectors. Additionally, we introduce $\pi^{*}$-index that clarifies the connection between its underlying concept and sequential coalitional games (see Sec. 4.3), and prove that this index satisfies the essential axioms of coalitional games (see App. C.9).
@@ C.2 Details of the algorithm
Figure H: Conceptual illustration of reward computation. Given a partially corrupted image, the detector outputs multiple proposals. The reward is computed by scoring each proposal against the target detection in terms of localization and classification agreement, and taking the maximum score.
C.3 Conceptual illustration of reward computation
Figure H illustrates the concept of the reward computation in Eq. (3). Given a partially corrupted image $x_{S}$, the detector outputs multiple proposals. For each proposal, the reward evaluates how well it matches the target detection $(L^{x},P^{x})$ in terms of both localization and classification agreement. The final reward is defined as the maximum score over all proposals.
C.4 Specific example of patch insertion
@@ C.5 Patch deletion
C.6 Specific example of patch deletion
We provide specific examples for cases with $r=1$ and $r=2$ in patch deletion described in App. C.5.
C.6.1 Case with r=1
@@ C.6.2 Case with r=2
C.7 Heat maps generation from identified patches
@@ C.8 Implementation details
Figure I: Comparison results of the two filling methods, zero filling and blurred pixel values, using (a) the insertion metric and (b) the deletion metric. The similarity score is computed using Eq. (3).
Figure J: The comparison of insertion and deletion metrics between VX-CODE and Shapley values estimated via Monte Carlo sampling.
Figure K: Comparison of heat maps generated by VX-CODE with patch insertion for the original and weight-randomized models.
@@
C.9 Axiomatic properties of the $\pi^{*}$-index
@@
The $\psi^{\mathrm{es}}$ measures the contribution of each player by the decrease in reward upon their removal. It is worth noting that the greedy patch deletion process described in App. C.5 can be interpreted as the player elimination setting in a sequential coalitional game. As shown in Eq. 62, at each step the algorithm selects the set of patches whose removal produces the largest decrease in the reward, which is equivalent to selecting the set with the largest full-context Shapley value. In this sense, the patch deletion process procedure induces a specific permutation $\hat{\pi}$ that defines a sequential coalitional game in the player elimination setting.
@@ Appendix D Additional experimental results
Figure L: Additional results comparing the visual explanations generated by each method for detections from DETR. In VX-CODE, we set $r=1$. (a) Heat maps generated by each method. (b) Patches identified from VX-CODE with patch insertion (left) and patch deletion (right). Each heat map corresponds to those shown in (a).
Figure M: Additional results comparing the visual explanations generated by each method for detections from Faster R-CNN. In VX-CODE, we set $r=1$. (a) Heat maps generated by each method. (b) Patches identified from VX-CODE with patch insertion (left) and patch deletion (right). Each heat map corresponds to those shown in (a).
Figure N: Additional results comparing the identified patches and generated heat maps by VX-CODE when changing $r$. (a) Results for patch insertion. (b) Results for patch deletion.
Figure O: (a) Example of synthetic data: A square marker is placed in the upper-right corner outside the bounding box. These data include annotations for the class and bounding box. (b) Example of data without annotations: These data lack annotations and are treated as background during training.
@@ D.1 Comparison of filling methods
As described in Sec. 5, we fill the masked regions with zeros in the proposed method. Here, we compare two filling methods, zero filling and blurred pixel values, using the insertion and deletion metrics. Figure I shows the results. We observe that both methods produce nearly identical results. These findings suggest that the choice of filling method has no significant impact on the proposed method.
@@ D.2 Comparison of Monte-Carlo based Shapley values
We compare the proposed method with Shapley values estimated via Monte Carlo sampling. Figure J shows the results of the insertion and deletion metrics, computed for 100 instances detected by DETR on MS-COCO. The number of Monte Carlo samples is set to 200, following prior studies, which already incurs a higher computational cost than our method with $r=1$. As shown in Fig. J, the proposed method with $r=1$ already outperforms the Shapley values estimated by Monte Carlo sampling.